54007GD:

A jar contains 0.25 liter of apple juice and 0.30 liter of grape juice. Melissa poured 0.75 liter of pineapple juice into the jar. She then drank 0.20 liter of the mixture. Part A: Write an expression to represent the total amount of juice left in the jar. Part B: Simplify the expression and identify which property is used in each step.

Brazyxkapalot:

the awsner is 1.10 u add the orignal jusice and the added juice and then u subtract the juice that has been gone and then thas how u get the awnser of how much is left it tht make sense. 0.25 + 0.30 + 0.75) - 0.20 = 1.30 - 0.20 = 1.10 left in jar

Starrysky6105:

(original amount of juice + added juice) - juice that was drank = juice left in the jar 0.25 + 0.3 + 0.75 - 0.2 = the answer would be 1.1 liter

Rozzy678:

nicoleonfleek (original amount of juice + added juice) - juice that was drank = juice left in the jar (0.25 + 0.30 + 0.75) - 0.20 = 1.30 - 0.20 = 1.10 left in jar

Rozzy678:

1.10